# Safe Planning in Unknown Environments using Conformalized Semantic Maps

### v1

@@ I Introduction
Autonomous robot navigation has received considerable research attention, leading to the development of numerous planners capable of designing trajectories that reach known goal regions while maintaining safe distances from obstacles [1, 2, 3]. Recent advances in AI-enabled robot vision and semantic mapping have created new opportunities to move beyond these well-studied geometric planning approaches toward semantic planning, where task completion requires reasoning not only about the geometric structure of the environment (e.g., obstacle locations) but also its semantic structure (e.g., object categories or terrain types) [4, 5].
@@
This paper focuses on semantic planning problems in environments populated with multiple objects and regions whose locations and semantic categories are initially unknown. The robot’s goal is to reach desired destinations while maintaining category-dependent safety distances from these semantic regions and objects. Recent planning algorithms tackle such tasks by leveraging AI-enabled perception systems (e.g., segmentation and object recognition algorithms) to collect geometric and semantic observations [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17]. These approaches typically either (i) employ deep learning methods to directly map perceptual inputs to robot actions (e.g., [15, 14, 17]) or (ii) leverage semantic SLAM techniques [18, 19, 20] to build semantic maps that are used by downstream planners (e.g., [6, 7, 21]). Despite impressive empirical performance, in these architectures the planning layer generally treats outputs from the perception or mapping modules as ground truth, ignoring their inherent uncertainty. As a result, these approaches lack mission completion guarantees. Attempts to model perceptual uncertainty within semantic and geometric planning have been investigated in [6, 7, 22] and [23, 24, 25, 26, 27], respectively. However, these methods often depend on unreliable proxies such as SLAM-derived confidence scores, or assume known sensor models and Gaussian noise—assumptions that may not hold in practice. Specifically: (i) SLAM-based uncertainty metrics often stem from fusing uncalibrated confidence scores from perception systems [28]; (ii) sensor models are often nonlinear and poorly characterized in real environments; and (iii) noise is often non-Gaussian or multi-modal.
@@
Figure 1: Illustration of the proposed planning framework for semantic reach-avoid tasks $\psi$. Given categorical and range measurements, a semantic map is constructed [19]. To account for map uncertainty caused by perceptual imperfections, conformal prediction is employed to generate sets of maps that contain the ground-truth environment with a user-defined probability. These sets are then used by a planner to generate uncertainty-aware paths that complete the assigned tasks with a user-defined probability $1-\alpha$.
@@
This paper addresses these limitations by developing a semantic planner for unknown environments that completes semantic reach-avoid tasks with a user-defined probability under perceptual uncertainty, while remaining agnostic to sensor models and noise characteristics. Robots are equipped with imperfect perception systems (e.g., cameras with semantic segmentation) that provide geometric and semantic observations. Our framework consists of three main components; see Fig. 1. First, we leverage existing semantic mapping algorithms that capture dense semantic and geometric environmental information and generate a multi-class semantic map of the environment [19]. Second, we quantify uncertainty of the generated semantic maps using Conformal Prediction (CP), a statistical, model- and distribution-free tool [29]. This produces a set of maps guaranteed to contain the ground truth with a user-defined probability. Third, we develop a semantic planner that balances exploitation and exploration. In exploitation mode, the planner selects the most conservative map from the CP set and generates paths using any search- or sampling-based planner [1, 2, 3]. When uncertainty is too high for a feasible solution to exist, the planner switches to exploration mode to reduce uncertainty (modeled by the size of the CP set) until a feasible path can be found. We show both theoretically and empirically that our method achieves desired mission completion rates. We also conduct experiments highlighting the exploitation–exploration trade-offs as the required success probability increases. Our comparative experiments also demonstrate that our approach outperforms baselines that either treat mapping outputs as ground truth or rely on uncertainty metrics provided by the mapping algorithm.
@@
Related Work: (i) Perception-based Planning: As discussed earlier, numerous planners for unknown environments have been proposed. However, these either neglect perceptual/mapping uncertainty and, thus, lack correctness guarantees [8, 9, 10, 11, 13, 14, 12, 15, 16, 17], or assume known sensor models and noise characteristics to design probabilistically correct paths [6, 7, 23, 24, 25, 26, 27]. In contrast, our planner achieves user-specified mission completion rates without assuming any knowledge of sensor models or noise. The closest work to ours is [30], which employs CP to quantify the uncertainty of object detectors and design probabilistically correct paths. However, [30] focuses solely on geometric planning and cannot straightforwardly handle combined geometric and semantic uncertainty. (ii) CP for Verified Autonomy: CP has recently been applied to quantify the uncertainty of various AI-enabled components to ensure verified autonomy; see the survey in [31]. Applications include trajectory prediction models [32, 33], large language models [34, 35], decision-making algorithms [36], image classifiers [37], diffusion dynamic models [38], and 3D human pose estimators [39]. Inspired by these works, we propose the first use of CP to quantify semantic map uncertainty, thereby enabling probabilistically correct semantic planning.
@@
Contributions: The contributions of this paper can be summarized as follows. First, we introduce the first semantic planner for unknown environments that achieves user-defined mission completion rates under perceptual uncertainty, while remaining agnostic to sensor models and noise characteristics. Second, we present the first application of CP to quantify the uncertainty of semantic maps generated by existing mapping algorithms, also in a model-agnostic manner. Third, we present experiments that validate the theoretical mission completion guarantees and show that our method consistently outperforms baselines in terms of mission success rates.
@@ II-C Perception System and Semantic Mapping
The robot is equipped with a sensor of range $r_{\text{max}}>0$ that provides information about the distances to, and semantic categories of, surrounding objects along a set of rays. We denote by $\mathcal{Z}_{t}:=\{z_{t,b}\}_{b}$ the set of measurements collected along ray $b$ at time $t$, where each measurement is given by $z_{t,b}=\{r_{t,b},y_{t,b}\}$, with $r_{t,b}\in\mathbb{dR}_{\geq 0}$ and $y_{t,b}\in{\mathcal{K}}$ denoting the range and categorical measurements, respectively. Such information can be obtained by processing RGB-D images using a semantic segmentation algorithm [40]; see Fig. 2.
These labeled range measurements can be fused online to build a semantic representation of $\Omega$ using existing semantic mapping algorithms. In particular, we adopt [19], which models the environment as a multi-class grid map $\mathbf{m}$} defined as a collection of cells $j\in{\mathcal{J}}=\{1,\dots,M\}$, each labeled with a category $m_{j}\in{\mathcal{K}}$ and represented by its center $c_{j}\in\Omega$; see Fig. 2. To account for perceptual noise, each cell $m_{j}$ maintains a probability mass function (PMF) $p_{t}(m_{j})=p(m_{j}|{\mathbf{x}}_{0:t},{\mathcal{Z}}_{0:t})$ over $\mathcal{K}$, conditioned on the measurements ${\mathcal{Z}}_{0:t}$ collected along the robot path ${\mathbf{x}}_{0:t}$. If at some $t$, no measurement corresponds to cell $j$, $p_{t}(m_{j})=p_{t-1}(m_{j})$. Accordingly, the distribution of map is defined as $p_{t}({\mathbf{m}})=\prod_{j\in{\mathcal{J}}}p_{t}(m_{j})$ [19]. For brevity, we denote by $\sigma$ a mapping algorithm that takes as input the distributions $p_{t-1}(m_{j})$ and the new measurement ${\mathcal{Z}}_{t}$ collected at ${\mathbf{x}}_{t}$, and outputs the updated distributions $p_{t}(m_{j})$ for all $j\in{\mathcal{J}}_{t}$, i.e., $p_{t}({\mathbf{m}})=\sigma(\Omega,p_{t-1}({\mathbf{m}}),{\mathcal{Z}}_{t})$, where ${\mathcal{J}}_{t}\subseteq{\mathcal{J}}$ is the set of indices of the grid cells that have been detected till time $t$; map update at $t=0$ occurs using a prior $p_{-1}({\mathbf{m}})$. Hereafter, when it is clear from the context, we refer to the random variable ${\mathbf{m}}_{t}$ without explicitly specifying its density $p_{t}({\mathbf{m}})$.
@@ II-D Problem Statement
We assume that mission scenarios $\xi=({\mathbf{x}}_{0},\Omega,\psi)$, determined by the initial robot state ${\mathbf{x}}_{0}$, an environment $\Omega$, and a semantic reach-avoid task $\psi$, are sampled from an unknown distribution $\mathcal{D}$}. While $\mathcal{D}$} is unknown, we assume that i.i.d. scenarios $\xi$ can be sampled from it. As discussed above the robot does not know $\Omega$; instead it has access to a prior map ${\mathbf{m}}_{-1}$ that can be updated online using the mapping algorithm $\sigma$. We aim to design a planning algorithm that, given a scenario sampled from $\mathcal{D}$}, can generate a robot path ${\mathbf{x}}_{0:H}$, for some $H>0$, that completes $\psi$ with a user-specified probability $1-\alpha$, given the sequence of maps ${\mathbf{m}}_{0:H}$ updated online. In summary, we address the following problem:
@@
Consider a mission scenario $\xi=({\mathbf{x}}_{0},\Omega,\psi)\sim\mathcal{D}$, where $\Omega$ is unknown but prior information ${\mathbf{m}}_{-1}$ is available. Given a mapping algorithm $\sigma$, design a planning algorithm that generates a robot path ${\mathbf{x}}_{0:H}$, for some $H>0$, ensuring that
$\mathbb{P}_{\xi\sim{\mathcal{D}}}({\mathbf{x}}_{0:H}\models\psi~|~{\mathbf{m}}_{0:H})\geq 1-\alpha,$ (2)
where and $1-\alpha\in(0,1)$ is a user-specified success rate.
@@
(i) The robot is equipped with an omnidirectional sensor of range $r_{\text{max}}$ generating $\mathcal{Z}_{t}$. (ii) The sensor range $r_{\text{max}}$ satisfies $r_{\text{max}}\geq d_{k}$ for all $k\in{\mathcal{K}}$ (iii) If a cell $j$ in the map is occupied by an object lying within the sensor’s field of view, a measurement associated with it will always be returned. Due to (iii), we have that if no measurements are obtained for a cell $j$ (lying within the sensor range), then it corresponds to free-space.
@@ III Safe Semantic Planning in Unknown Environments
A natural approach to Problem II.1 is to plan directly over the distribution of the map ${\mathbf{m}}_{t}$, as in [7, 6]. However, in practice, the sensor models used to update the map distribution are inaccurate, so the PMFs $p_{t}(m_{j})$ do not represent true probabilities. As a result, planning directly over ${\mathbf{m}}_{t}$ may compromise the probabilistic correctness of the generated paths. To overcome this limitation, we propose a new planner that makes no assumptions about the sensor model and noise characteristics. In Section III-A, we employ conformal prediction (CP) to ‘calibrate’ ${\mathbf{m}}_{t}$ by computing prediction sets for each cell $j\in{\mathcal{J}}_{t}$ that contain the true label with user-defined confidence. These sets are then used in Section III-B to design paths ${\mathbf{x}}_{0:H}$ that solve Problem II.1.
@@ III-A Conformalizing Semantic Maps
As discussed earlier, the distribution of ${\mathbf{m}}_{t}$ does not represent true probabilities compromising the correctness of paths designed only using ${\mathbf{m}}_{t}$. To address this, for any test scenario $\xi_{\text{test}}\sim{\mathcal{D}}$ associated with an unknown environment $\Omega$, we aim to design sets ${\mathcal{C}}_{j}({\mathbf{m}}_{t})\subseteq{\mathcal{K}}$ of labels for each cell $j\in{\mathcal{J}}_{t}$ detected along a path ${\mathbf{x}}_{0:H}$ and time step $t\in\{0,\dots,H\}$ satisfying the following property:
$\mathbb{P}_{\xi\sim{\mathcal{D}}}\left(C_{0:H}\mid{\mathbf{m}}_{0:H}\right)\geq 1-\alpha,$ (3)
where $C_{0:H}=\bigwedge_{\forall j\in{\mathcal{J}}_{t},t\in[0:H]}(k_{j}^{\text{gt}}\in{\mathcal{C}}_{j}({\mathbf{m}}_{t}))$ and $k_{j}^{\text{gt}}\in{\mathcal{K}}$ is the true label of the region occupying cell $j$. In words, (3) guarantees that, with probability at least $1-\alpha$, the sets ${\mathcal{C}}_{j}({\mathbf{m}}_{t})$ contain the true labels $k_{j}^{\text{gt}}$ for all cells $j\in{\mathcal{J}}_{t}$ in ${\mathbf{m}}_{t}$, all time steps $t\in\{0,\dots,H\}$, and any robot path ${\mathbf{x}}_{0:H}$ followed to construct ${\mathbf{m}}_{0:H}$.
To construct these sets for $\xi_{\text{test}}$, we employ CP to quantify uncertainty in the outputs of the mapping algorithm $\sigma$. Applying CP requires (i) a set of calibration mission scenarios $\{\xi_{i}=({\mathbf{x}}_{0,i},\Omega_{i},\psi_{i})\}_{i=1}^{D}$ drawn from $\mathcal{D}$} with known ground-truth labels $k_{j}^{\text{gt}}$ for all $j\in{\mathcal{J}}$, and (ii) a non-conformity score (NCS) that captures the error of $\sigma$, which generates maps ${\mathbf{m}}_{0:H}$ along any path ${\mathbf{x}}_{0:H}$. It is crucial that the NCS accounts for mapping errors along arbitrary robot paths, rather than a single fixed path, since otherwise (3) would hold only for that path. Thus, we adopt the following NCS, which is path-agnostic (and, consequently, task-agnostic) and captures the worst-case mapping error, associated with the calibration mission scenario $\xi_{i}$:
@@
where $\mathcal{P}$} is the set of all possible finite paths in the calibration environment $\Omega_{i}$. In (4), the term $p_{t}(m_{j}=k_{j}^{\text{gt}})$ denotes the probability that cell $j$ has the ground-truth label according to the map ${\mathbf{m}}_{t}$. Consequently, $1-p_{t}(m_{j}=k_{j}^{\text{gt}})$ models the mapping error at cell $j$ after collecting measurements along a specific path ${\mathbf{x}}_{0:t}$. The maximum operator in (4) computes the worst-case error, i.e., the largest error across all cells $j\in{\mathcal{J}}_{t}$, all time steps $t\in[0,H]$, and all paths ${\mathbf{x}}_{0:H}\in{\mathcal{P}}$. [^1] Computing (4) exactly requires evaluating the maximum over an infinite number of paths, which is impractical. As in [30], in practice the NCS is approximated using a finite set of paths $\mathcal{P}$}; see Section V-C for details. We compute the NCSs $\{s_{i}\}_{i=1}^{D}$ for all calibration scenarios and determine their $\frac{\lceil(D+1)(1-\alpha)\rceil}{D}$-quantile, denoted by $\hat{s}$.
For an unseen/test mission scenario $\xi_{\text{test}}\sim{\mathcal{D}}$ with unknown true labels for cells $j$, at each step $t$ we compute the following prediction set of labels for all cells $j\in{\mathcal{J}}_{t}$ (see Alg. 1):

### v2

@@ I Introduction
Extensive research on autonomous robot navigation has resulted in numerous planners capable of designing trajectories that reach known goal regions while avoiding obstacles [a-star, 12]. Recent advances in AI-enabled robot vision and semantic mapping have created new opportunities to move beyond these well-studied geometric planning approaches toward semantic planning, where task completion requires reasoning not only about the geometric structure of the environment (e.g., obstacle locations) but also its semantic structure (e.g., object categories or terrain types) [8, 19].
@@
This paper focuses on semantic planning problems in environments populated with multiple objects and regions whose locations and semantic categories are initially unknown. The robot’s goal is to reach desired destinations while maintaining category-dependent safety distances from these semantic regions and objects. Recent planning algorithms tackle such tasks by leveraging AI-enabled perception systems (e.g., object recognition algorithms) to collect geometric and semantic observations [11, fu2016optimal, 28, 15, 17, 21]. These approaches typically either (i) employ deep learning methods to directly map perceptual inputs to robot actions (e.g., [7, 15, 21]) or (ii) leverage semantic SLAM techniques [2, ssmi_temp, 27] to build semantic maps that are used by downstream planners (e.g., [11, fu2016optimal, 20]). Despite impressive empirical performance, in these architectures the planning layer generally treats outputs from the perception or mapping modules as ground truth, ignoring their inherent uncertainty. As a result, these approaches lack mission completion guarantees. Attempts to model perceptual uncertainty within semantic and geometric planning have been investigated in [11, fu2016optimal, 31] and [24, 3, 18, 9], respectively. However, these methods often depend on unreliable proxies such as SLAM-derived confidence scores that do not represent true probabilities [16], or assume known sensor models and Gaussian noise—assumptions that may not hold in practice.
@@
Figure 1: Illustration of the proposed planning framework for semantic reach-avoid tasks $\psi$. Given categorical and range measurements, a semantic map is constructed [ssmi_temp]. To account for map uncertainty caused by perceptual imperfections, conformal prediction is employed to generate sets of maps that contain the ground-truth environment with a user-defined probability. These sets are then used by a planner to generate uncertainty-aware paths that complete the assigned tasks with a user-defined probability $1-\alpha$.
@@
This paper addresses these limitations by developing a semantic planner for unknown environments that completes semantic reach-avoid tasks with a user-defined probability under perceptual uncertainty, while remaining agnostic to sensor models and noise characteristics. Robots are equipped with imperfect perception systems (e.g., cameras with semantic segmentation) that provide geometric and semantic observations. Our framework consists of three main components; see Fig. 1. First, we leverage existing semantic mapping algorithms that capture dense semantic and geometric environmental information and generate a multi-class semantic map of the environment [ssmi_temp]. Second, we quantify uncertainty of the generated semantic maps using Conformal Prediction (CP), a statistical, model- and distribution-free tool [CPbase]. This produces a set of maps guaranteed to contain the ground truth with a user-defined probability. Third, we develop a semantic planner that balances exploitation and exploration. In exploitation mode, the planner selects the most conservative map from the CP set and generates paths using any search- or sampling-based planner [a-star, 12]. When uncertainty is too high for a feasible solution to exist, the planner switches to exploration mode to reduce uncertainty (modeled by the size of the CP set) until a feasible path can be found. We show both theoretically and empirically that our method achieves desired mission completion rates. We also conduct experiments highlighting the exploitation–exploration trade-offs as the required success probability increases. Our comparative experiments also demonstrate that our approach outperforms baselines that either treat mapping outputs as ground truth or rely on uncertainty metrics provided by the mapping algorithm.
@@
Related Work: (i) Perception-based Planning: Numerous planners for unknown environments have been proposed. However, these either neglect perceptual/mapping uncertainty and, thus, lack correctness guarantees [22, 28, 15, 7, 17, 21], or assume known sensor models and noise characteristics to design probabilistically correct paths [11, fu2016optimal, 24, 3, 18, 9]. In contrast, our planner achieves user-specified mission completion rates without assuming any knowledge of sensor models or noise. The closest work to ours is [PwC], which employs CP to quantify the uncertainty of object detectors and design probabilistically correct paths. However, [PwC] focuses solely on geometric planning and cannot straightforwardly handle combined geometric and semantic uncertainty. (ii) CP for Verified Autonomy: CP has been applied to quantify uncertainty in various AI-enabled components (e.g., perception, motion prediction, and LLMs) [14, 25, 5, 23, 30, 13, 26]. Inspired by these works, we propose the first use of CP to quantify semantic map uncertainty enabling probabilistically correct semantic planning.
@@
Contributions: First, we introduce the first semantic planner for unknown environments that achieves user-defined mission completion rates under perceptual uncertainty, while remaining agnostic to sensor models and noise characteristics. Second, we present the first application of CP to quantify the uncertainty of semantic maps generated by existing mapping algorithms, also in a model-agnostic manner. Third, we present experiments that validate the theoretical mission completion guarantees and show that our method consistently outperforms baselines in terms of mission success rates.
@@ II-C Perception System and Semantic Mapping
The robot is equipped with a sensor of range $r_{\text{max}}>0$ that provides information about the distances to, and semantic categories of, surrounding objects along a set of rays. We denote by $\mathcal{Z}_{t}:=\{z_{t,b}\}_{b}$ the set of measurements collected along ray $b$ at time $t$, where each measurement is given by $z_{t,b}=\{r_{t,b},y_{t,b}\}$, with $r_{t,b}\in\mathbb{dR}_{\geq 0}$ and $y_{t,b}\in{\mathcal{K}}$ denoting the range and categorical measurements, respectively.
These labeled range measurements, obtained by processing RGB-D images, can be fused online to build a semantic representation of $\Omega$ using semantic mapping algorithms. We adopt [ssmi_temp], which models the environment as a multi-class grid map $\mathbf{m}$} defined as a collection of cells $j\in{\mathcal{J}}=\{1,\dots,M\}$, each labeled with a category $m_{j}\in{\mathcal{K}}$ and represented by its center $c_{j}\in\Omega$; see Fig. 2. To account for perceptual noise, each cell $m_{j}$ maintains a probability mass function (PMF) $p_{t}(m_{j})=p(m_{j}|{\mathbf{x}}_{0:t},{\mathcal{Z}}_{0:t})$ over $\mathcal{K}$, conditioned on the measurements ${\mathcal{Z}}_{0:t}$ collected along the robot path ${\mathbf{x}}_{0:t}$. If at some $t$, no measurement corresponds to cell $j$, $p_{t}(m_{j})=p_{t-1}(m_{j})$. Accordingly, the distribution of map is defined as $p_{t}({\mathbf{m}})=\prod_{j\in{\mathcal{J}}}p_{t}(m_{j})$. For brevity, we denote by $\sigma$ a mapping algorithm that takes as input the distributions $p_{t-1}(m_{j})$ and the new measurement ${\mathcal{Z}}_{t}$ collected at ${\mathbf{x}}_{t}$, and outputs the updated distributions $p_{t}(m_{j}),\forall j\in{\mathcal{J}}_{t}$, i.e., $p_{t}({\mathbf{m}})=\sigma(\Omega,p_{t-1}({\mathbf{m}}),{\mathcal{Z}}_{t})$, where ${\mathcal{J}}_{t}\subseteq{\mathcal{J}}$ is the set of indices of the grid cells that have been detected till time $t$; map update at $t=0$ occurs using a prior $p_{-1}({\mathbf{m}})$. Hereafter, when it is clear from the context, we refer to the random variable ${\mathbf{m}}_{t}$ without explicitly specifying its density $p_{t}({\mathbf{m}})$.
@@ II-D Problem Statement
Consider any user-defined initial robot state ${\mathbf{x}}_{0}$ and semantic reach-avoid task $\psi$ for an environment $\Omega$. We assume that $\Omega$ is sampled from an unknown distribution $\mathcal{D}$}, from which a finite set of i.i.d environments can be sampled. The robot does not know $\Omega$; instead it has access to a prior map ${\mathbf{m}}_{-1}$ that can be updated online using the mapping algorithm $\sigma$. We aim to design a planning algorithm that, given an environment sampled from $\mathcal{D}$}, can generate a robot path ${\mathbf{x}}_{0:H}$, for some $H>0$, that completes $\psi$ with a user-specified probability $1-\alpha$, given the sequence of maps ${\mathbf{m}}_{0:H}$ updated online. Formally, we address the following problem:
@@
Consider an unknown environment $\Omega\sim{\mathcal{D}}$ with prior information ${\mathbf{m}}_{-1}$, and any initial robot state ${\mathbf{x}}_{0}$ and semantic reach-avoid task $\psi$ defined over $\Omega$. Given a mapping algorithm $\sigma$, design a planning algorithm that generates a robot path ${\mathbf{x}}_{0:H}$, for some $H>0$, ensuring that
$\mathbb{P}_{{\Omega\sim{\mathcal{D}}}}({\mathbf{x}}_{0:H}\models\psi~|~{\mathbf{m}}_{0:H})\geq 1-\alpha,$ (2)
where $1-\alpha\in(0,1)$ is a user-specified success rate.
@@
(i) The robot is equipped with an omnidirectional sensor of range $r_{\text{max}}$ generating $\mathcal{Z}_{t}$. (ii) The sensor range $r_{\text{max}}$ satisfies $r_{\text{max}}\geq d_{k}$ for all $k\in{\mathcal{K}}$ (iii) Any cell $j$ in the map within the sensor’s field of view always has a measurement associated with it. Due to (iii), if no measurements are obtained for cell $j$ (lying within the sensor range), then it corresponds to free-space.
@@ III Safe Semantic Planning in Unknown Environments
A natural approach to Problem II.1 is to plan directly over the distribution of the map ${\mathbf{m}}_{t}$, as in [fu2016optimal, 11]. However, in practice, the sensor models used to update the map distribution are inaccurate, so the PMFs $p_{t}(m_{j})$ do not represent true probabilities. As a result, planning directly over ${\mathbf{m}}_{t}$ may compromise the probabilistic correctness of the generated paths. To overcome this limitation, we propose a new planner that makes no assumptions about the sensor model and noise characteristics. In Section III-A, we employ conformal prediction (CP) to ‘calibrate’ ${\mathbf{m}}_{t}$ by computing prediction sets for each cell $j\in{\mathcal{J}}_{t}$ that contain the true label with user-defined confidence. These sets are then used in Section III-B to design paths ${\mathbf{x}}_{0:H}$ that solve Problem II.1.
@@ III-A Conformalizing Semantic Maps
As discussed earlier, the distribution of ${\mathbf{m}}_{t}$ does not represent true probabilities, compromising the correctness of paths designed only using ${\mathbf{m}}_{t}$. To address this, for any test environment $\Omega_{\text{test}}\sim{\mathcal{D}}$, we aim to design sets ${\mathcal{C}}_{j}({\mathbf{m}}_{t})\subseteq{\mathcal{K}}$ of labels for each cell $j\in{\mathcal{J}}_{t}$ detected along a path ${\mathbf{x}}_{0:H}$ and time step $t\in\{0,\dots,H\}$ satisfying the following property:
$\mathbb{P}_{{\Omega\sim{\mathcal{D}}}}\left(C_{0:H}\mid{\mathbf{m}}_{0:H}\right)\geq 1-\alpha,$ (3)
where $C_{0:H}=\bigwedge_{\forall j\in{\mathcal{J}}_{t},t\in[0:H]}(k_{j}^{\text{gt}}\in{\mathcal{C}}_{j}({\mathbf{m}}_{t}))$ and $k_{j}^{\text{gt}}\in{\mathcal{K}}$ is the true label of the region occupying cell $j$. In words, (3) guarantees the inclusion of the true labels $k_{j}^{\text{gt}}$ in the sets ${\mathcal{C}}_{j}({\mathbf{m}}_{t})$$\forall j\in{\mathcal{J}}_{t}$ in ${\mathbf{m}}_{t}$, and $\forall t\in\{0,\dots,H\}$, and any robot path ${\mathbf{x}}_{0:H}$ followed to construct ${\mathbf{m}}_{0:H}$.
To construct these sets for $\Omega_{\text{test}}$}, we employ CP to quantify uncertainty in the outputs of the mapping algorithm $\sigma$. Applying CP requires (i) a set of calibration environments ${\{\Omega_{i}\}_{i=1}^{D}}\sim{\mathcal{D}}$ with known ground-truth labels $k_{j}^{\text{gt}}$ for all $j\in{\mathcal{J}}$, and (ii) a non-conformity score (NCS) that captures the error of $\sigma$, which generates maps ${\mathbf{m}}_{0:H}$ along any path ${\mathbf{x}}_{0:H}$. It is crucial that the NCS accounts for mapping errors along arbitrary robot paths, rather than a single fixed path, since otherwise (3) would hold only for that path. Thus, we adopt the following NCS, which is path (and, consequently, task)-agnostic and captures the worst-case mapping error, associated with the calibration environment $\Omega_{i}$:
@@
where $\mathcal{P}$} is the set of all possible finite paths in $\Omega_{i}$. In (4), the term $p_{t}(m_{j}=k_{j}^{\text{gt}})$ denotes the probability that cell $j$ has the ground-truth label according to the map ${\mathbf{m}}_{t}$. Consequently, $1-p_{t}(m_{j}=k_{j}^{\text{gt}})$ models the mapping error at cell $j$ after collecting measurements along a specific path ${\mathbf{x}}_{0:t}$. The maximum operator in (4) computes the worst-case error, i.e., the largest error across all cells $j\in{\mathcal{J}}_{t}$, all time steps $t\in[0,H]$, and all paths ${\mathbf{x}}_{0:H}\in{\mathcal{P}}$. [^1] Computing (4) exactly requires evaluating the maximum over an infinite number of paths, which is impractical. As in [PwC], in practice the NCS is approximated using a finite set of paths $\mathcal{P}$}; see Section V-A for details. We compute the NCSs $\{s_{i}\}_{i=1}^{D}$ for all calibration environments and determine their $\frac{\lceil(D+1)(1-\alpha)\rceil}{D}$-quantile, denoted by $\hat{s}$.
For an unseen test environment $\Omega_{\text{test}}\sim{\mathcal{D}}$ with unknown true labels for cells $j$, at each step $t$ we compute the following prediction set of labels $\forall j\in{\mathcal{J}}_{t}$ (see Alg. 1):